Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-08-02

Observing: Morgan Fraser, Tom Reynolds
Support Team: Seppo Mattila
TAT: Lluís Galbany, Stefan Taubenberger


Weather conditions

Start of night:  Clear, seeing ~1"
Middle of night: Clear, seeing ~1"
End of night: Clear, seeing <1"

------------------------------------------------

Calibrations

Evening: iV flats
Morning: grz flats

------------------------------------------------

23:03    Standard, LTT7379 (low counts due to slit problem, see below)
23:13    AT2016eld (classification): Somehow a nearby bright object was placed in the slit. See below
23:37    End of twilight
            Image Analysis
23:53    SN2016AJ (follow-up): Observed 16x100 ri
01:12    AT2016eld (class):  Observed, but target was not in slit, as slit had moved from expected position. This was why we had failed acquisition earlier
01:35    SN2016blz (follow-up): Observed Gr11+16 2700s, but target was not necessarily in the slit, see above
03:24    AT2016emu (class): Attempted, could not see target in acquisition image
            Realised slit error here, and corrected.
03:29    AT2016els (class): Observed, probable SLSNe (!)
04:28    ASASSN15oz (follow-up): Observed 2700s, gr13, 1.0" slit
05:44    OGLE16aaa (follow-up): Observed 3600 Gr13, 1.0" slit
06:30    OGLE15xl (follow-up): Observed 9x200 gri
08:29    SMT-DWF16bc (class): Young Ia
09:08    ASASSN15og (follow-up): Observed 1x2700s & 1x1800s Gr11
09:59    Beginning of Twilight
10:37    Standard, Feige110
 

Comments